# Scaling Laws for Upcycling Mixture-of-Experts Language Models

### v1

@@ A.6 Upcycled training’s configuration
We initialize the router weights from a normal distribution with zero mean and variance of $2/5d_{\rm model}$ [Le Scao et al., 2022, Chowdhery et al., 2023, Zeng et al., 2022] (the same initialization is used for from-scratch trainings).
Regarding the learning rate (LR) of upcycled training, there are several choices: using the LR at the end of dense training with a constant LR schedule, i.e., treating upcycled training as a kind of fine tuning; using the LR of dense/MoE training [Komatsuzaki et al., 2022, He et al., 2024]. We consider these three choices without retuning the LR. With the other training settings are set to be the same as the one used in training the dense models, including the use of the WSD LR schedule, we find that using the MoE LR leads to better performance. See Figure 8.
@@
Let us further comment on alternative upcycling methods that could potentially further accelerate training. Adding some form of noises to the dense MLP weights, or modifying partially the weights would intuitively help upcycled training generalize faster. Our preliminary experiments however did not see any advantages of doing so. Note that this observation is also consistent with previous negative reports [Komatsuzaki et al., 2022, Wei et al., 2024, Muennighoff et al., 2024]. Henceforth, we simply copy the weights directly from the dense model to perform upcycled training.
@@ A.8 Ablation of dataset repetition for upcycling
In our experiments, we have used the same dataset for training both dense and upcycled models, following practices of [He et al., 2024]. As an ablation, we split our dataset to two non-overlapping portions, and perform the two-stage training with the distinct datasets. In Figure 10, we find that the difference in performance is very small. This also aligns with previous investigation on data repetition, where it is shown that there is a little difference in performance up to 4 times of repetition [Muennighoff et al., 2023].
@@ A.9 Evaluation with standard benchmarks and comparison with other existing models
We compare the performance of our trained 1B models against existing models with similar sizes, Pythia [Biderman et al., 2023] and TinyLlama [Zhang et al., 2024b], based on standard natural language processing benchmarks, ARC [Clark et al., 2018], lambada [Paperno et al., 2016], logiqa [Liu et al., 2020], piqa [Bisk et al., 2020], sciq [Welbl et al., 2017], and winogrande [Sakaguchi et al., 2021].
@@
Table 7: Benchmarks’ performance comparison across models. Reported scores are accuracies (normalized by byte length whenever applicable). The first two columns are scores of existing models. The last two columns are evaluation results of models trained in this work. The upcycled 8x1B model is upcycled from the 1B model. Our models are evaluated with the LM Evaluation Harness v0.4.0 library [Gao et al., 2024].
@@ A.10 Estimated Total GPU hours
The estimate is as follows. We calculate the FLOPs for training the dense, MoE, upcycled MoE models with maximum iterations using the $6ND$ approximation, ignoring the additional FLOPs required to continued pretrain models with shorter iterations (as we can reuse the intermediate checkpoints). We further assume that the per-second TFLOPs of the GPU is 400, and is the same for both dense and MoE models.  [^6] https://github.com/NVIDIA/Megatron-LM/blob/main/megatron/core/transformer/moe/README.md. Note that MoE requires more GPU memory to store its total model parameters, inducing overhead that may slow down training. We obtain,
@@ Appendix B Generalization across other datasets
We show that the scaling behavior studied in the main text generalizes across datasets in use. Aside from Slimpajama [together2023redpajama, cerebras2023slimpajama, Shen et al., 2023] used in the main text, we test with two additional datasets: Japanese portion of the CommonCrawl corpus, and the Stack code dataset [Kocetkov et al., 2022]. We use a tokenizer of enlarged vocabulary size of 102,400 when training with the former dataset.  [^7] https://huggingface.co/sbintuitions/sarashina2-70b We upcycle a 0.1B model with the training configuration set to the same as those used for the experiments presented in the main text. The results are presented in Figure 11, where the scaling behaves similarly to those presented in Figure 4, suggesting that our scaling law for dataset are independent of dataset in use. We however do not further study scaling law for model configuration and the joint scaling law due to computational constraints.
@@
Appendix C More results on scaling law for dataset sizes
@@ C.3 Bilinear form
Appendix D More Results on the Joint Scaling Law
D.1 Validating Requirement 3
@@
D.2 Another functional form satisfying Requirements 3 and 4
@@
D.3 Why fitting scaling laws separately?
@@
Appendix E More implication of the Joint Scaling Law
E.1 From-scratch training vs Upcycling
@@
E.2 More on compute optimality of upcycling
@@
E.3 Derivation of compute-optimal formula

### v2

@@ A.6 Upcycled training’s configuration
We initialize the router weights from a normal distribution with zero mean and variance of $2/5d_{\rm model}$ (Le Scao et al., 2022; Chowdhery et al., 2023; Zeng et al., 2022) (the same initialization is used for from-scratch trainings).
Regarding the learning rate (LR) of upcycled training, there are several choices: using the LR at the end of dense training with a constant LR schedule, i.e., treating upcycled training as a kind of fine tuning; using the LR of dense/MoE training (Komatsuzaki et al., ; He et al., 2024). We consider these three choices without retuning the LR. With the other training settings are set to be the same as the one used in training the dense models, including the use of the WSD LR schedule, we find that using the MoE LR leads to better performance. See Figure 8.
@@
Let us further comment on alternative upcycling methods (parameter reinitialization) that could potentially further accelerate training. Adding some form of noises to the dense MLP weights, or modifying partially the weights would intuitively help upcycled training generalize faster. Our preliminary experiments (adding Gaussian noise to the expert weights of the upcycled model; partially randomizing the MLP weights; and substituting the MLP weights with their low-rank couterparts randomly to encourage expert divergence) however did not see any advantages of doing so. Note that this observation is also consistent with previous negative reports (Komatsuzaki et al., ; Wei et al., 2024; Muennighoff et al., 2024). Henceforth, we simply copy the weights directly from the dense model to perform upcycled training.
@@ A.8 Ablation of dataset repetition for upcycling
In our experiments, we have used the same dataset for training both dense and upcycled models, following practices of (He et al., 2024). As an ablation, we split our dataset to two non-overlapping portions, and perform the two-stage training with the distinct datasets. In Figure 10, we find that the difference in performance is very small. This also aligns with previous investigation on data repetition, where it is shown that there is a little difference in performance up to 4 times of repetition (Muennighoff et al., 2023).
@@ A.9 Evaluation with standard benchmarks and comparison with other existing models
We compare the performance of our trained 1B models against existing models with similar sizes, Pythia (Biderman et al., 2023) and TinyLlama (Zhang et al., 2024), based on standard natural language processing benchmarks, ARC (Clark et al., 2018), lambada (Paperno et al., 2016), logiqa (Liu et al., 2021), piqa (Bisk et al., 2020), sciq (Welbl et al., 2017), and winogrande (Sakaguchi et al., 2021).
@@
Table 7: Benchmarks’ performance comparison across models. Reported scores are accuracies (normalized by byte length whenever applicable). The first two columns are scores of existing models. The last two columns are evaluation results of models trained in this work. The upcycled 8x1B model is upcycled from the 1B model. Our models are evaluated with the LM Evaluation Harness v0.4.0 library (Gao et al., 2024).
@@ A.10 Estimated Total GPU hours
The estimate is as follows. We calculate the FLOPs for training the dense, MoE, upcycled MoE models with maximum iterations using the $6ND$ approximation, ignoring the additional FLOPs required to continued pretrain models with shorter iterations (as we can reuse the intermediate checkpoints). We further assume that the per-second TFLOPs of the GPU is 400, and is the same for both dense and MoE models  [^7] https://github.com/NVIDIA/Megatron-LM/blob/main/megatron/core/transformer/moe/README.md. Note that MoE requires more GPU memory to store its total model parameters, inducing overhead that may slow down training.. We obtain,
@@ Appendix B Generalization across Other Datasets
We show that the scaling behavior studied in the main text generalizes across datasets in use. Aside from Slimpajama (Computer, 2023; Soboleva et al., 2023; Shen et al., 2023) used in the main text, we test with two additional datasets: Japanese portion of the CommonCrawl corpus, and the Stack code dataset (Kocetkov et al., ). We use a tokenizer of enlarged vocabulary size of 102,400 when training with the former dataset  [^8] https://huggingface.co/sbintuitions/sarashina2-70b. We upcycle a 0.1B model with the training configuration set to the same as those used for the experiments presented in the main text. The results are presented in Figure 11, where the scaling behaves similarly to those presented in Figure 4, suggesting that our scaling law for dataset is independent of dataset in use. We however do not further study scaling law for model configuration and the joint scaling law due to computational constraints.
We further observe:
  1. 1.
    The Japanese dataset has higher validation loss (harder task), while the code dataset results in lower loss, making it a relatively easier task in terms of cross-entropy loss.
  2. 2.
    Interestingly, the Japanese dataset is harder to saturate with increasing $D_{1}$, meaning upcycled training remains effective. In contrast, the code dataset saturates more quickly, making upcycling less beneficial.
The more fine-grained question, e.g., how mixtures of pretraining data impact respective downstream performances, is beyond the scope of our current work, and we leave it for future study.
@@
Appendix C More Results on the Scaling Law for Dataset Sizes
@@ C.3 Bilinear form
Appendix D More Results on the Scaling Law for Model Configuration
D.1 Fitting the Scaling Law for Model Configuration
Recall that we wish to understand the cross-entropy loss as a function of sparsity, defined as $P=N_{\rm total}/N_{2}$ and the number of active parameters, $N_{2}$. Starting from the power-law ansatz, we require that the loss satisfies
$L(P,N_{2})=L_{P}(N_{2})=AN_{2}^{-\beta_{1}}+E$
and
$L(P,N_{2})=L_{N_{2}}(P)=AP^{-\beta_{2}}+E$
This is reasonable as $P$, when fixing $N_{2}$, is the total number of model parameters, which we expect to satisfy the power-law ansatz. Analogously, $N_{2}$ corresponds to the number of dense model parameters, which should satisfy the power-law ansatz as well.
We then consider, as before, both additive and multiplicative functional forms, with and without interaction, satisfying the above requirements, and evaluate them using leave-one-out RMS error:
Multiplicative:
$L(P,N_{2})$ $=AP^{-\beta_{1}}(1+N_{2})^{-\beta_{2}+\beta_{3}\log P}+E$
$\approx AP^{-\beta_{1}}N_{2}^{-\beta_{2}+\beta_{3}\log P}+E$ (17)
Additive:
$L(P,N_{2})$ $=AP^{-\beta_{1}}+F(1+N_{2})^{-\beta_{2}+\beta_{3}\log P}+E$
$\approx AP^{-\beta_{1}}+FN_{2}^{-\beta_{2}+\beta_{3}\log P}+E$ (18)
We obtain RMS errors of 0.0414, 0.0351, 0.0341 and 0.0322 for multiplicative (with interaction), multiplicative (without interaction), additive (with interaction), and additive (without interaction) cases, respectively. That is, the additional form without interaction functional form provides the best fit to the data.
D.2 Alternative Visualization of the Scaling Law for Model Configuration
We show the 2D slices (with fixed sparsity) of the model configuration scatter plot as in the left panel of Figure 1, in Figure 13.
Figure 13: Scaling behavior with respect to model configuration. Shown are plots of the cross-entropy loss with respect to active parameter $N_{2}$ fixing sparsity $P$.
Appendix E More Results on the Joint Scaling Law
E.1 Validating Requirement 3
@@
E.2 Another functional form satisfying Requirements 3 and 4
@@
E.3 Why fitting scaling laws separately?
@@
Appendix F More Implication of the Joint Scaling Law
F.1 From-scratch training vs Upcycling
@@
F.2 More on compute optimality of upcycling
@@
F.3 Derivation of compute-optimal formula